Determine the average value of the form factor of a half wave rectifier for the alternating current represent by
a. i = 20sinwt
b. i = 30sinwt
1
Expert's answer
2021-09-04T01:01:13-0400

a)

"\text{For half wave siniusoidal waveform}\\
i_{avg}=\dfrac{i_{max}}{2}\\
\text{Form factor}=1.57\\
i=20\sin\omega{t}\\
i_{max}=20A\\
i_{avg}=\dfrac{20}{2}=10A\\" b)

"i=30\sin\omega{t}\\
i_{max}=30A\\
i_{avg}=\dfrac{30}{2}=15A\\"